Differentiate cos x w.r.t. ex.
Please answer with proper steps

Grade:11

2 Answers

Arun
4 years ago
Dear student
 
first differentiate cos x w.r.t. x
 
 d(cos x) /dx = – sin x
 
differentiate e^x w.r.t. x
 
d(e^x)/dx = e^x
 
now d(cos x) / d(e^x) = – sin x / (e^x)
Pintu Chaudhary
4 years ago
Hello, friend, try this method
        let, y=cosx
               \small \tfrac{dy}{de^{x}}=\tfrac{d cosx}{de^{x}}
 
                      \small =\tfrac{d cosx}{dx}\times\tfrac{dx}{de^{x}}
                      =\tfrac{d cosx}{dx}\times\tfrac{1}{\tfrac{de^{x}}{dx}}
                       = -sinx \times \frac{1}{^{e^{x}}}
                       =\tfrac{-sinx}{e^{x}}  Answer. 
      If you get it then fine but if you don't then message for the steps which you don't. Thanks...